OpenStudy (anonymous):

(2x^3-3x-10)/(x-2) using synthetic division

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

A=2x^2-4x+5 b=2x^2+4x+5 c=2x^2+4x-5 d=x^2-4x-5+1/(x-2)

OpenStudy (anonymous):

those are the only options

OpenStudy (anonymous):

im confused let me see

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

your welcome